CLISP(1) CLISP(1) NAME clisp−CommonLisplanguageinterpreterandcompiler SYNOPSIS clisp[-h][-mmemsize][-sstacksize][-Mmemfile][-q][-iinitfile...][-c[-l]lispfile ...][-xexpression] DESCRIPTION Invokesthecommonlispinterpreterandcompiler.Invokedwithoutarguments,executesareadeval-printloop,inwhichexpressionsareinturnreadfromstandardinput,evaluatedbythelisp interpreter,andtheirresultsoutputtostandardoutput.Invokedwith−c,thespecifiedlispfiles arecompiledtoabytecodethatcanbeexecutedmoreefficiently. OPTIONS -hDisplaysahelpmessageonhowtouseclisp. -mmemsize Setstheamountofmemoryclisptriestograbonstartup.Theamountmaybegivenas nnnnnnn(measuredinbytes),nnnnKornnnnKB(measuredinkilobytes)ornMor nMB(measuredinmegabytes).Defaultis2.5megabytes.Theargumentisconstrained between100KBand16MB.--Thisversionofclispallocatesmemorydynamically. memsizeisessentiallyignored. -sstacksize Setsthesizeofthestackclispallocatesforitself.Thesyntaxisthesameasformemsize. Defaultisoneeighthofmemsize.Theargumentisconstrainedbetween40KBand8 MB. -Mmemfile Specifiestheinitialmemoryimage.Thismustbeamemorydumpproducedbythe saveinitmemfunction. -qQuiet:clispdisplaysnobanneratstartupandnogood-byemessagewhenquitting. -iinitfile... Specifiesinitializationfilestobeloadedatstartup.Theseshouldbelispfiles(sourceor compiled). -clispfile... Compilesthespecifiedlispfilestobytecode.Thecompiledfilescanthenbeloadedinstead ofthesourcestogainefficiency. -lAbytecodelistingofthefilesbeingcompiledwillbeproduced.Usefulonlyfordebugging purposes. -xexpressions Executesaseriesofarbitraryexpressionsinsteadofaread-eval-printloop.Thevaluesof theexpressionswillbeoutputtostandardoutput.Duetotheargumentprocessingdone bytheshell,theexpressionsmustbeenclosedinsinglequotes,anddoublequotesand backslashesmustbeprecededbybackslashes. REFERENCE Thelanguageimplementedconformsto GuyL.SteeleJr.:CommonLisp-TheLanguage. DigitalPress.1stedition1984,465pages. ("CLtL1"forshort) andtotheolderpartsof GuyL.SteeleJr.:CommonLisp-TheLanguage. DigitalPress.2ndedition1990,1032pages. ("CLtL2"forshort) 20December1993 1 CLISP(1) CLISP(1) USE helptogetsomeon-linehelp. (aproposname) liststhesymbolsrelatingtoname. (exit)or(quit)or(bye) toquitclisp. EOF(Ctrl-Z) toleavethecurrentread-eval-printloop. FILES lisp.exe mainexecutable lispinit.mem initialmemoryimage config.lsp site-dependentconfiguration *.lsplispsource *.faslispcode,compiledbyclisp *.liblispsourcelibraryinformation,generatedandusedbytheclispcompiler ENVIRONMENT COMSPEC(DOS,OS/2implementationsonly) isusedtofindthecommandinterpretercalledbythefunctionshell. SEEALSO cmulisp(1),emacs(1). BUGS Thefunctioninspectisnotimplemented. OnlyveryfewextensionsfromCLtL2aresupported. Noon-linedocumentationbeyondaproposanddescribeisavailable. PROJECTS Writingon-linedocumentation. Buildingaforeignfunctioninterface(abilitytocallCcodedirectly). Writeinspect. Enhancethecompilersuchthatitcaninlinelocalfunctions. Specifyaportablesetofwindowandgraphicsoperations. AUTHORS BrunoHaible<haible@ma2s2.mathematik.uni-karlsruhe.de>andMichaelStoll. 20December1993 2